The World of 3G

mclab

October 21, 2003

The World of 3G

1

2

[width=5.25in,height=3.05in]seminar0401.eps

Figure 1:

1. 3

[width=4.96in,height=3.69in]seminar04021.eps

Figure 2:

4

(a) 5

Technical Information The CDMA2000 standard is evolving to continually support new services in a standard 1.25 MHz carrier. The first phase of CDMA2000, or CDMA2000 1X will deliver average data rates of 144 kbps. Phase two, labeled CDMA2000 1xEV, will provide for data rates greater than 2Mbps.

CDMA2000 1X The IS-2000 standard (CDMA2000 1X) was completed early this year and published by the Telecommunications Industry Association (TIA). 1X offers approximately twice the voice capacity of cdmaOne, average data rates of 144 kbps, backward compatibility with cdmaOne networks, and many other performance improvements. 1X refers to CDMA2000 implementation within existing spectrum allocations for cdmaOne - 1.25 MHz carriers. The technical term is derived from N =1 (i.e. use of same 1.25 MHz carrier as in cdmaOne) and the 1x means one times 1.25 MHz.

[width=4.17in,height=1.33in]seminar04031.eps

CDMA2000 1X can be implemented in existing spectrum or in new spectrum allocations. A CDMA2000 1X network will also introduce simultaneous voice and data services, low latency data support and other performance improvements. The backward compatibility with cdmaOne provided by Ericsson’s CDMA2000 solution, further ensures investment protection. CDMA2000 1xEV The evolution of CDMA2000 beyond 1X is now labeled CDMA2000 1xEV. 1xEV will be divided into two steps: 1xEV-DO and 1xEV-DV. 1xEV-DO stands for 1X Evolution Data Only. 1xEV-DV stands for 1X Evolution Data and Voice. Both 1xEV evolution steps provide for advanced services in CDMA2000 using a standard 1.25 MHz carrier. Evolution with CDMA2000 will therefore continue to be backwards compatible with today’s networks and forwards compatible with each evolution option! 1xEV-DO will be available for CDMA2000 operators during 2002, and will provide for higher data rates on 1X systems. 1xEV-DO will require a separate carrier for data, but this carrier will be able to hand-off to a 1X carrier if simultaneous voice and data services are needed. By allocating a separate carrier for data, operators will be able to deliver peak rates in excess of 2 Mbps (best effort) to their data customers. 1xEV-DV solutions will be available approximately one and a half to two years after 1xEV-DO. IXEV-DV will bring data and voice services for CDMA2000 back into one carrier. A 1xEV-DV carrier will provide not only high speed data and voice simultaneously, but will also be capable of delivering real-time packet services. CDMA2000 Packet Core Network The standards for a CDMA packet core network are being developed by the TR45.6 working group of the TIA. These standards are being developed by using existing standards from the IETF (Internet Engineering Task Force) on Mobile IP. 3GPP2 is also busy defining the evolution of a CDMA2000 network to All-IP. The CDMA2000 PCN will be the first step in this evolution. To provide secure and efficient transport for wireless data, Ericsson has introduced a data delivery mechanism using PCN (Packet Core Network). The PCN network is comprised of the PDSN (Packet Data Serving Node) and AAA (Authentication, Authorization and Accounting). The HA (Home Agent) can be added to provide Mobile IP based packet data services. Ericsson is leading the industry in the development of a standardized PCN with open interfaces for CDMA Operators.

[width=6.51in,height=3.06in]seminar04041.eps

cdmaOne Evolution Options Ericsson provides customers with the freedom to choose how and when to evolve to a 3G network solution. Ericsson’s cdmaOne and CDMA2000 products are designed to protect current and future network investment. With Ericsson’s CDMA2000 air interface and core solution, operators can double the voice capacity of existing cdmaOne systems and introduce 3G features and services without significant capital investment in infrastructure and spectrum.